angle 1 and 2 are complementary. If angle 1 was eight degrees less and angle 2 was fifty-seven degrees more, then they would be of equal measure. What are the measures of the two angles?

Answer by stanbon(75887)

angle 1 and 2 are complementary. If angle 1 was eight degrees less and angle 2 was fifty-seven degrees more, then they would be of equal measure. What are the measures of the two angles?
-----------
Let angle 1 be "x"; then angle 2 is "90-x"
EQUATION:
x-8 = 90-x + 57
2x = 155
x = 72.5 degrees
90-x = 17.5 degrees
===============
Cheers,
Stan H.